OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

are u a baby this is such a simple question? well it is -2/5

ganeshie8 (ganeshie8):

AX+BY+C = 0 slope of a line in standard form : -A/B

OpenStudy (lgbasallote):

first put it in the form y= mx + b start by subtracting 2x from both sides \[5y = -2x - 10\] now we want y only so divide both sides by 5 \[y = \frac{-2x - 10}{5}\] now break it down \[y = -\frac{2x}{5} - \frac{10}{5}\] \[y = -\frac{2}{5}x - 2\] the slope is the coefficient of x

OpenStudy (anonymous):

AX+BY+C = 0 u can use this formula